Write an equation in Slope-intercept form for the line that is perpendicular to the line y = 2x - 5 that passes through the given point (4,-2). Show your work.
Answer by algebrahouse.com(1659)

Perpendicular lines have slopes which are negative reciprocals.

In the line y = 2x - 5,
the slope is 2

The perpendicular slope is -1/2

(4,-2) , m = -1/2 {point and slope}
y - y1 = m(x - x1) {point-slope form}
y - (-2) = (-1/2)(x - 4) {substituted point and slope into point-slope form}
y + 2 = (-1/2)x + 2 {used distributive property}
y = (-1/2)x {subtracted 2 from each side}

For more help from me, visit: www.algebrahouse.com